OpenStudy (shamallamadingdong):

Help plz Which is the equation of the line that contains points (8, 10) and (–4, 2)?

OpenStudy (shamallamadingdong):

\[y-8=\frac{ 3 }{ 2 }(x-10)\] \[y-10=\frac{ 2 }{ 3 }(x-8)\] \[y-8=\frac{ 2 }{ 3 }(x-10)\] \[y-10=\frac{ 3 }{ 2 }(x-8)\]

OpenStudy (shamallamadingdong):

@Lihtkah

OpenStudy (anonymous):

oh my. hold on. just a sec or so.

OpenStudy (anonymous):

You need to expand out the options, then sub in the values to see if they work.

OpenStudy (anonymous):

The gradient of the curve is 2/3

OpenStudy (shamallamadingdong):

@pooja195